# Feature Space Analysis by Guided Diffusion Model

### v1

@@ Abstract
One of the key issues in Deep Neural Networks (DNNs) is the black-box nature of their internal feature extraction process. Targeting vision-related domains, this paper focuses on analysing the feature space of a DNN by proposing a decoder that can generate images whose features are guaranteed to closely match a user-specified feature. Owing to this guarantee that is missed in past studies, our decoder allows us to evidence which of various attributes in an image are encoded into a feature by the DNN, by generating images whose features are in proximity to that feature. Our decoder is implemented as a guided diffusion model that guides the reverse image generation of a pre-trained diffusion model to minimise the Euclidean distance between the feature of a clean image estimated at each step and the user-specified feature. One practical advantage of our decoder is that it can analyse feature spaces of different DNNs with no additional training and run on a single COTS GPU. The experimental results targeting CLIP’s image encoder, ResNet-50 and vision transformer demonstrate that images generated by our decoder have features remarkably similar to the user-specified ones and reveal valuable insights into these DNNs’ feature spaces.
Index Terms:
@@
Feature space, Diffusion model, Guidance, Multimodal embedding space.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Deep learning is often referred to as representation learning or feature learning that extracts useful features (data representations) through statistical analysis on a large amount of data [1, 2]. In vision-related domains, this is typically demonstrated by impressive image classification performances achieved through an output layer that performs simple linear classification on top of extracted features [3, 4]. In addition, regardless of whether a deep neural network (DNN) is pre-trained or not, features extracted by it can be used as a source to carry out various tasks. For example, cross-modal retrieval is implemented via similarity search on features (embeddings) extracted by encoders trained on different modalities [5, 6], image/video captioning is performed by passing features extracted by an encoder to a decoder [7], and image editing is done using features extracted from some inputs (e.g., image, text prompt, sketch, etc.) as conditions [8, 9].
@@
Despite the above-mentioned extensive exploration of utilising features extracted by DNNs, very limited attention has been given to understanding those features. In addition, it is now known that overparameterised DNNs, which memorise atypical examples and outliers/mislabeled examples, offer an improved generalisation power rather than overfitting [10, 11]. This motivates us to investigate what kind of feature space is built by such a DNN. Also, visual-semantic embeddings are learned so that an image and its relevant text caption are encoded into close points (i.e., embeddings) in a common feature space, but they are not the same [5, 6]. This evokes our interest in what kind of image (or caption) is characterised by an embedding located between those of the relevant image and caption.
@@
To analyse the feature space of a DNN, this paper addresses the development of a decoder that generates an image whose feature closely matches a user-specified one. This decoder allows us to visually reveal which of various attributes in an image are encoded into a feature extracted by the DNN, by generating images whose features are in proximity to this feature. Furthermore, it could be possible to generate, for example, a “hard negative” image whose feature is difficult to distinguish from the one extracted from a semantically dissimilar image [5], a “rare” image whose feature lies far from the feature of every image in a dataset, and a “memorised” image for which a small perturbation of its feature leads to a different prediction [12]. Therefore, the decoder developed in this paper opens up many possibilities for analysing and improving the feature space of the DNN (mining the above-mentioned images is our future work).
@@
    Our decoder generates images whose features are ensured to closely match user-specified ones, which enables us to perform rigorous analysis of a DNN’s feature space based on the short Euclidean distances between those features. To our best knowledge, we are the first to propose this kind of rigorous feature space analysis. In particular, the experiments targeting CLIP’s image encoder, ResNet-50 and Vision Transformer (ViT) reveal valuable insights into their feature spaces, such as little sensitivity of CLIP’s image encoder to an object’s anatomical structure, loss of detailed information by legacy ResNet-50, ViT’s excessive focus on the main object in an image, and weak image-text association in CLIP’s feature space.
@@
    Our decoder is general and training-free. By general, it can be used to analyse the feature space of any DNN that encodes an image into a feature. By training-free, our decoder needs no additional training, that is, its feature space analysis can be performed as long as one implements a feature extraction process by a DNN.
@@ II Related Work
Many guidance approaches have been developed to provide different controls over the image generation of a diffusion model in order to devise various applications such as text-to-image [18, 19, 20, 22, 25, 24, 26, 27], image inpainting [20, 26, 27], object/style/layout manipulation [21, 23, 26, 27], composition between images [23] and image segmentation [21]. Regarding loss functions used in existing guided diffusion models, the most basic ones like classifier guidance [18], classifier-free guidance [19] and CLIP guidance [20] measure a probability or score indicating how relevant a text prompt is to a noisy image generated at each step. These loss functions are then extended to examine whether each of subject tokens in a prompt is attended to by some patches in a generated image [22], whether a generated image satisfies propositional statements deduced from a prompt [25] and whether a generated image has a plausible depth map [24]. In addition, various loss functions are proposed to assess different aspects such as an object’s position, size, shape and appearance [23], an object’s pose and position, personal identity and painting style [26, 27] and the distribution of colors and the one of labels in a generated segmentation image [21].
@@
To analyse the feature space of a DNN, some researchers have investigated decoders that invert a feature back into an image from which this feature is likely to be extracted. Teterwak et al. adopts the framework of conditional GAN to train a decoder (generator) in which the operation of batch normalisation is modulated based on a given feature (to be precise logit vector) [13]. Ramesh et al. construct a decoder as a diffusion model that generates an image by combining a given feature with the embedding of each step and also deriving four extra textual tokens of context from it [14]. Bordes et al. build a decoder as a diffusion model by developing conditional batch normalisation layers that take a given feature as conditioning [15]. As inferred from the above-mentioned studies, in principle, one can devise a decoder as any conditional generative model that generates an image, which is likely to be encoded into a user-specified feature.
However, this approach involves two crucial problems: First, just using a user-specified feature as conditioning lacks checking how close the feature of a generated image is to it. The experiment in Section IV-B shows that although images generated by a conditional diffusion model in [15] are plausible, their features are distant from the corresponding specified features. Second, training generative models on large-scale datasets requires very expensive computational costs. In particular, a diffusion model should be employed as a decoder because of its capability to generate high-quality images, but training it is prohibitively expensive. For instance, training the diffusion model in [18] consumes 150150-$1000$ V100 GPU days to produce high-quality images [28]. Moreover, training stable diffusion [17] demands over $24$ days on $256$ A100 GPUs [29]. Therefore, it is infeasible for ordinary researchers to train a diffusion-model decoder on a large-scale dataset. Therefore, we employ a guided diffusion model that is based on a pre-trained diffusion model and needs no additional training.
@@
III Our Guided-diffusion-model-based Decoder for Feature Space Analysis
@@ III-B Details of Our Decoder
3:  Set the number of self-recurrence iterations $K$ based on the early step emphasis
@@
The overall object layout of a generated image is primarily established during early reverse steps [16]. Thus, our decoder’s image generation fails if this layout is inadequate. To address this, we design an early step emphasis strategy that prioritises early reverse steps. Specifically, for the first $T^{\prime}$ reverse steps (i.e., $t\geq T-T^{\prime}$ in Algorithm 1) the number of self-recurrence iterations $K$ is set to a large value while using a much smaller $K$ at the remaining steps. In our case, $K$ is set to $1000$ and $8$ at the first $T^{\prime}=5$ reverse steps and the others, respectively. Appendix C provides an ablation study to empirically show the effectiveness of the early step emphasis strategy.
Figure 2: Distributions of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ obtained at the first self-recurrence (i.e., $k=0$) iteration of the first seven reverse steps (i.e., t=999t=999-$993$th steps) when the target feature extractor $f$ is defined as CLIP’s image encoder (a), ResNet-50 (b) and ViT-H/14 (c). Here, $\boldsymbol{f}_{s}$ is specified as the feature extracted from the top-left actual image in Fig. 3 (please see Section IV for more details). In each visualised distribution, the dotted indicates the mean and the solid lines represent three times standard deviation for (a) and (b) and twice standard deviation for (c).
Stable diffusion’s decoder produces an image $\hat{\boldsymbol{x}}_{t}$ as a tensor of shape $C\times H\times W$ (in our case $C=3$ and $H=W=512$) where pixels can take both negative and positive values. The ideal way to precisely compute $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ is to scale values in $\hat{\boldsymbol{x}}_{t}$ and save $\hat{\boldsymbol{x}}_{t}$ into an actual image file, which is then loaded and encoded into $f\left(\hat{\boldsymbol{x}}_{t}\right)$ through $f$. However, this is infeasible because the computation graph is broken when $\hat{\boldsymbol{x}}_{t}$ is saved into an external image file. Thus, to simulate image save on a GPU’s VRAM, a virtual image save module is implemented so that $\hat{\boldsymbol{x}}_{t}$ is scaled and rounded to make pixels take discrete values between 0 and $255$  [^5] Due to the small difference in image rescaling, $f\left(\hat{\boldsymbol{x}}_{t}\right)$ obtained via the visual image save differs marginally from $f\left(\hat{\boldsymbol{x}}_{t}\right)$ obtained by loading an actual image file. In the case where $f$ is defined as CLIP’s image encoder to produce $1024$-dimensional feature vectors, their squared Euclidean distance is about $0.011$, resulting in only minor semantic variations in generated images.. Here, straight through estimator [34] is used to approximate the backpropagation over the round function of $\hat{\boldsymbol{x}}_{t}$ as its identify function.
@@
In particular, as depicted by solid lines in Fig. 2, $\nabla_{thres}$ is set to three times the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ when $f$ is CLIP’s image encoder or ResNet-50. When using ViT-H/14 as $f$, $\nabla_{thres}$ is set to twice the standard deviation due to the broad spread of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ as shown in Fig. 2 (c). Finally, $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ fixed by the above-mentioned gradient normalisation and clipping is weighted by $w_{g}$ to compute $\epsilon_{\theta}^{\prime}\left(\boldsymbol{z}_{t},t\right)$ in Eq. (3). All the experiments in the next section are done by fixing $w_{g}$ to $4$. Please see Appendixes B and D for results obtained by different values of $w_{g}$ and $\nabla_{thres}$, respectively.
IV Experimental Results
We aim to evaluate our decoder from two perspectives: The first is the evaluation as a decoder, specifically whether our decoder can successfully generate images whose features closely match a user-specified feature $\boldsymbol{f}_{s}$. The second perspective is the evaluation as a feature space analyser, namely whether generated images provide meaningful visual cues about the feature space of the target feature extractor $f$. One key issue in designing experiments is determining how to define $\boldsymbol{f}_{s}$ that serves as the goal—that is, the feature of a generated image should be as similar to $\boldsymbol{f}_{s}$ as possible. We define $\boldsymbol{f}_{s}$ as the feature that is extracted from an actual image via $f$. This also facilitates the evaluation of the second perspective because we can interpret the adequacy of visual contents in a generated image by referring to the actual image.
@@ IV Experimental Results
By setting $f$ to each of the three feature extractors described above, $\boldsymbol{f}_{s}$ is extracted from every one of the $15$ actual images shown in the leftmost column in Figs. 3 and 4. These images are selected from the dataset used in [26, 27] (the upper three images in Fig. 3), MSCOCO dataset [35] (the five images from the bottom of Fig. 3 and the uppermost image in Fig. 4), and ImageNet dataset [36] (the six images except the uppermost one in Fig. 4). The reason for the small number of image generation attempts is our decoder’s computation time as discussed in Section IV-D. Another reason is the limitation of computational resources, as all the experiments were conducted using only one NVIDIA GeForce RTX 3090 with 2424GB VRAM and one NVIDIA RTX 4000 Ada with 2020GB VRAM. However, as seen from the leftmost column in Figs. 3 and 4, the $15$ images encompass a variety of objects (e.g., person, animal, vehicle and food) presented with different backgrounds, so we believe that they are enough for evaluating the overall performance of our decoder.
The hyper-parameters specific to our decoder ($K$, $T^{\prime}$ and $w_{g}$) remain fixed across all the experiments, except for $\nabla_{thres}$ that is set differently when $f$ is ViT or the others. Of course, better results might be obtained if finer hyper-parameter tuning were conducted for each experiment. Please see Section III-B for the exact hyper-parameter values. The configuration of the stable diffusion in our decoder is the same as the one in [26, 27].
IV-A Results Targeting CLIP’s Image Encoder
Figs. 3 and 4 show images generated by our decoder when using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor. For each row, $\boldsymbol{f}_{s}$ is extracted as the feature of the leftmost actual image and our decoder’s task is to generate an image whose feature is as similar to $\boldsymbol{f}_{s}$ as possible in terms of their squared Euclidean distance. Considering the randomness in generated images, our decoder is run three times. For each run, the following two images are presented: The first is the “finally generated image” that is obtained at the end of the image generation process, and the second is the “best image” whose feature yields the shortest squared Euclidean distance to $\boldsymbol{f}_{s}$ during the image generation process. The number under each generated image shows the squared Euclidean distance between its feature and $\boldsymbol{f}_{s}$.

### v2

@@ Abstract
One of the key issues in Deep Neural Networks (DNNs) is the black-box nature of their internal feature extraction process. Targeting vision-related domains, this paper focuses on analysing the feature space of a DNN by proposing a decoder that can generate images whose features are guaranteed to closely match a user-specified feature. Owing to this guarantee that is missed in past studies, our decoder allows us to evidence which of various image attributes are encoded into the user-specified feature. Our decoder is implemented as a guided diffusion model that guides the reverse image generation of a pre-trained diffusion model to minimise the Euclidean distance between the feature of a clean image estimated at each step and the user-specified feature. One practical advantage of our decoder is that it can analyse feature spaces of different DNNs with no additional training and run on a single COTS GPU. The experimental results targeting CLIP’s image encoder, ResNet-50 and vision transformer demonstrate that images generated by our decoder have features remarkably similar to the user-specified ones and reveal valuable insights into these DNNs’ feature spaces.
keywords:
@@
Feature space analysis , Diffusion model , Guidance , Multimodal embedding space.
[…]
[1]organization=Department of Information Systems Design, Doshisha University,addressline=1-3 Tatara Miyakodani, city=Kyotanabe, postcode=610-0394, state=Kyoto, country=Japan
1 Introduction
@@
Deep learning is often referred to as representation learning or feature learning that extracts useful features (data representations) through statistical analysis on a large amount of data [1, 2]. In vision-related domains, this is typically demonstrated by impressive image classification performances achieved through an output layer that performs simple linear classification on top of extracted features [3, 4]. In addition, features extracted by a Deep Neural Network (DNN) can be used as a source to carry out various tasks. For example, cross-modal retrieval is implemented via similarity search on features (embeddings) extracted by encoders trained on different modalities [5, 6], image/video captioning is performed by passing features extracted by an encoder to a decoder [7], and image editing is done using features extracted from some inputs (e.g., image, text prompt, sketch, etc.) as conditions [8, 9].
@@
Despite the above-mentioned extensive exploration of utilising features extracted by DNNs, very limited attention has been given to understanding those features. It is now known that overparameterised DNNs, which memorise atypical examples and outliers/mislabeled examples, offer an improved generalisation power rather than overfitting [10, 11]. This motivates us to investigate what kind of feature space is built by such a DNN. Also, visual-semantic embeddings are learned so that an image and its relevant text caption are encoded into close points (i.e., embeddings) in a common feature space, but they are not the same [5, 6]. This evokes our interest in what kind of image (or caption) is characterised by an embedding located between those of the relevant image and caption.
@@
To analyse the feature space of a DNN, this paper addresses the development of a decoder that generates an image whose feature closely matches a user-specified one. This decoder allows us to visually reveal which of various attributes in an image are encoded into a feature extracted by the DNN, by generating images whose features are in proximity to this feature. Furthermore, it could be possible to generate, for example, a “hard negative” image whose feature is difficult to distinguish from the one extracted from a semantically different image [5], a “rare” image whose feature lies far from the feature of every image in a dataset, and a “memorised” image for which a small perturbation of its feature leads to a different prediction [12]. Therefore, the decoder developed in this paper opens up many possibilities for analysing and improving the feature space of the DNN (mining the above-mentioned images is our future work).
@@
    Our decoder generates images whose features are ensured to closely match user-specified ones, which enables us to perform rigorous analysis of the DNN’s feature space based on the short Euclidean distances between those features. To our best knowledge, we are the first to propose this kind of rigorous feature space analysis. In particular, the experiments targeting CLIP’s image encoder [6], ResNet-50 [3] and Vision Transformer (ViT) [4] reveal valuable insights into their feature spaces, such as little sensitivity of CLIP’s image encoder to an object’s anatomical structure, loss of detailed information by legacy ResNet-50, ViT’s excessive focus on the main object in an image, and weak image-text association in CLIP’s feature space.
@@
    Our decoder is general and training-free. By general, it can be used to analyse the feature space of any DNN that encodes an image into a feature. By training-free, our decoder needs no additional training, that is, feature space analysis can be performed as long as one implements the feature extraction process of a DNN.
@@ 2 Related Work
Many guidance approaches have been developed to provide different controls over the image generation of a diffusion model in order to devise various applications such as text-to-image [18, 19, 20, 22, 24, 25, 26, 27], image inpainting [20, 26, 27], object/style/layout manipulation [21, 23, 26, 27], composition between images [23] and image segmentation [21]. Regarding loss functions used in existing guided diffusion models, the most basic ones like classifier guidance [18], classifier-free guidance [19] and CLIP guidance [20] measure a probability or score indicating how relevant a text prompt is to a noisy image generated at each step. These loss functions are then extended to examine whether each of subject tokens in a prompt is attended to by some patches in a generated image [22], whether a generated image has a plausible depth map [24] and whether a generated image satisfies propositional statements deduced from a prompt [25]. In addition, various loss functions are proposed to assess different aspects such as an object’s position, size, shape and appearance [23], an object’s pose and position, personal identity and painting style [26, 27] and the distribution of colors and the one of labels in a generated segmentation image [21].
@@
To analyse the feature space of a DNN, some researchers have investigated decoders that invert a feature back into an image from which this feature is likely to be extracted. Teterwak et al. adopts the framework of conditional GAN to train a decoder (generator) in which the operation of batch normalisation is modulated based on a given feature (to be precise logit vector) [13]. Ramesh et al. construct a decoder as a diffusion model that generates an image by combining a given feature with the embedding of each step and also deriving four extra textual tokens of context from it [14]. Bordes et al. build a decoder as a diffusion model by developing conditional batch normalisation layers that take a given feature as conditioning [15]. As inferred from the above-mentioned studies, in principle, one can devise a decoder as any conditional generative model that is conditioned on a user-specified feature to generate an image, which is likely to be encoded into this feature.
However, this approach involves two crucial problems: First, just using a user-specified feature as conditioning lacks checking how close the feature of a generated image is to it. The experiment in Section 4.2 shows that although images generated by a conditional diffusion model in [15] are plausible, their features are distant from the corresponding specified features. Second, training generative models on large-scale datasets requires very expensive computational costs. In particular, a diffusion model should be employed as a decoder because of its capability to generate high-quality images, but training it is prohibitively expensive. For instance, training the diffusion model in [18] consumes 150150-$1000$ V100 GPU days to produce high-quality images [28]. Moreover, training stable diffusion [17] demands over $24$ days on $256$ A100 GPUs [29]. Therefore, it is infeasible for ordinary researchers to train a diffusion-model decoder on a large-scale dataset. We develop a guided diffusion model that not only explicitly examines the closeness between the feature of a generated image and a user-specified one, but also leverages a pre-trained diffusion model with no additional training.
@@
3 Our Guided-diffusion-model-based Decoder for Feature Space Analysis
@@ 3.2 Details of Our Decoder
3:  Set the number of self-recurrence iterations $K$ based on early step emphasis strategy
@@
The overall object layout of a generated image is primarily established during early reverse steps [16]. Thus, our decoder’s image generation fails if this layout is inadequate. To address this, we design an early step emphasis strategy that prioritises early reverse steps. Specifically, for the first $T^{\prime}$ reverse steps (i.e., $t\geq T-T^{\prime}$ in Algorithm 1) the number of self-recurrence iterations $K$ is set to a large value while using a much smaller $K$ at the remaining steps. In our case, $K$ is set to $1000$ and $8$ at the first $T^{\prime}=5$ reverse steps and the others, respectively. Part C in the supplementary materials provides an ablation study to empirically show the effectiveness of the early step emphasis strategy.
Figure 2: Distributions of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ obtained at the first self-recurrence iteration (i.e., $k=0$) of the first seven reverse steps (i.e., t=999t=999-$993$th steps) when the target feature extractor $f$ is defined as CLIP’s image encoder (a), ResNet-50 (b) and ViT-H/14 (c). Here, $\boldsymbol{f}_{s}$ is specified as the feature extracted from the top-left actual image in Fig. 3 (please see Section 4 for more details). In each visualised distribution, the dotted line indicates the mean and the solid lines represent three times the standard deviation for (a) and (b) and twice the standard deviation for (c).
Stable diffusion’s decoder produces an image $\hat{\boldsymbol{x}}_{t}$ as a tensor of shape $C\times H\times W$ (in our case $C=3$ and $H=W=512$) where pixels can take both negative and positive values. The ideal way to precisely compute $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ is to scale values in $\hat{\boldsymbol{x}}_{t}$ and save $\hat{\boldsymbol{x}}_{t}$ into an actual image file, which is then loaded and encoded into $f\left(\hat{\boldsymbol{x}}_{t}\right)$ through $f$. However, this is infeasible because the computation graph is broken when $\hat{\boldsymbol{x}}_{t}$ is saved into an external image file. Thus, to simulate image save on a GPU’s VRAM, a virtual image save module is implemented so that $\hat{\boldsymbol{x}}_{t}$ is scaled and rounded to make pixels take discrete values between 0 and $255$  [^5] Due to the small difference in image rescaling, $f\left(\hat{\boldsymbol{x}}_{t}\right)$ obtained via the visual image save module differs marginally from $f\left(\hat{\boldsymbol{x}}_{t}\right)$ obtained by loading an actual image file. In the case where $f$ is defined as CLIP’s image encoder to produce $1024$-dimensional feature vectors, their squared Euclidean distance is about $0.011$, resulting in only minor semantic variations in generated images.. Here, straight through estimator [34] is used to approximate the backpropagation over the round function of $\hat{\boldsymbol{x}}_{t}$ as its identify function.
@@
In particular, as depicted by solid lines in Fig. 2, $\nabla_{thres}$ is set to three times the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ when $f$ is CLIP’s image encoder or ResNet-50. When using ViT-H/14 as $f$, $\nabla_{thres}$ is set to twice the standard deviation due to the broad spread of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ as shown in Fig. 2 (c). Finally, $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ fixed by the above-mentioned gradient normalisation and clipping is weighted by $w_{g}$ to compute $\epsilon_{\theta}^{\prime}\left(\boldsymbol{z}_{t},t\right)$ in Eq. (3). All the experiments in the next section are done by fixing $w_{g}$ to $4$. Please see Parts B and D in the supplementary materials for results obtained by different values of $w_{g}$ and $\nabla_{thres}$, respectively.
4 Experimental Results
We aim to evaluate our decoder from two perspectives: The first is the evaluation as a decoder, specifically whether our decoder can successfully generate an image whose feature closely matches a user-specified feature $\boldsymbol{f}_{s}$. The second perspective is the evaluation as a feature space analyser, namely whether generated images provide meaningful visual cues about the feature space of the target feature extractor $f$. To design experiments that enable evaluations from these two perspectives, we define $\boldsymbol{f}_{s}$ as the feature that is extracted from an actual image via $f$. This is because we can interpret the adequacy of visual contents in a generated image by referring to the actual image.
@@ 4 Experimental Results
By setting $f$ to each of these three feature extractors, $\boldsymbol{f}_{s}$ is extracted from every one of the $15$ actual images shown in the leftmost column in Figs. 3 and 4. These images are selected from the dataset used in [26, 27] (the upper three images in Fig. 3), MSCOCO dataset [35] (the five images from the bottom of Fig. 3 and the uppermost image in Fig. 4), and ImageNet dataset [36] (the six images except the uppermost one in Fig. 4). The reason for the small number of image generation attempts is our decoder’s computation time as discussed in Section 4.4. Another reason is the limitation of our computational resources, as all the experiments were conducted using only one NVIDIA GeForce RTX 3090 with 2424GB VRAM and one NVIDIA RTX 4000 Ada with 2020GB VRAM. However, as seen from the leftmost column in Figs. 3 and 4, the $15$ images encompass a variety of objects (e.g., person, animal, vehicle and food) presented with different backgrounds, so we believe that they are enough for evaluating the overall performance of our decoder.
The hyper-parameters specific to our decoder ($K$, $T^{\prime}$ and $w_{g}$) remain fixed across all the experiments, except for $\nabla_{thres}$ that is set differently when $f$ is ViT or the others. Of course, better results might be obtained if finer hyper-parameter tuning were conducted for each experiment. Please see Section 3.2 for the exact hyper-parameter values. The configuration of stable diffusion in our decoder is the same as the one in [26, 27].
4.1 Results Targeting CLIP’s Image Encoder
Figs. 3 and 4 show images generated by our decoder when using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor. For each row, $\boldsymbol{f}_{s}$ is extracted as the feature of the leftmost actual image. Considering the randomness in generated images, our decoder is run three times. For each run, the following two images are presented: The first is the “finally generated image” that is obtained at the end of the image generation process, and the second is the “best image” whose feature yields the shortest squared Euclidean distance to $\boldsymbol{f}_{s}$ during the image generation process. The number under each generated image shows the squared Euclidean distance between its feature and $\boldsymbol{f}_{s}$.